Outcome of the CADE13 Business Meeting

David A. Plaisted

The outcome of the vote at the CADE 13 business meeting was that the trustees’ proposal passed, with some significant modifications. The first modification is that when the trustees cancel a motion passed at a business meeting, they are required to explain this in the AAR Newsletter. The second modification is that all elections and modifications to the bylaws will now be given to the entire AAR membership to vote on, and not just to the attendees at a business meeting. This is a significant improvement, and will give future democratic amendments a much better chance for approval. To pass, a constitutional amendment will require 30 percent of the members to vote, and of these, 2/3 must vote for approval.

The main drawback with the trustees’ proposal is that it is only partially democratic. That is, the upcoming program chairs and the secretary-treasurer have a vote among the trustees, even though these individuals were not elected and (probably) never served as CADE program chairs. This makes 3 or more non-elected trustees as well as six elected ones. My proposal would have specified that only elected individuals can be trustees. So we have the curious situation that the attendees at the business meeting voted to give themselves less of a voice in the running of CADE than they would have had under my proposal. I think that the main issues got lost in the technical complexities of the proposals at the business meeting.

Still, this is a tremendous step forward for CADE. I thank the trustees and the supporters of my proposal for what was accomplished, and hope that CADE will continue to become more democratic.